# Localized Schrödinger Bridge Sampler

### v1

@@ Abstract.
We consider the generative problem of sampling from an unknown distribution for which only a sufficiently large number of training samples are available. In this paper, we build on previous work combining Schrödinger bridges and Langevin dynamics. A key bottleneck of this approach is the exponential dependence of the required training samples on the dimension, $d$, of the ambient state space. We propose a localization strategy which exploits conditional independence of conditional expectation values. Localization thus replaces a single high-dimensional Schrödinger bridge problem by $d$ low-dimensional Schrödinger bridge problems over the available training samples. As for the original approach, the localized sampler is stable and geometric ergodic. The sampler also naturally extends to conditional sampling and to Bayesian inference. We demonstrate the performance of our proposed scheme through experiments on a Gaussian problem with increasing dimensions and on a stochastic subgrid-scale parametrization conditional sampling problem.
Keywords: generative modeling, Langevin dynamics, Schrödinger bridges, conditional independence, localization, Bayesian inference, conditional sampling, multi-scale closure
@@ 1. Introduction
In this paper, we consider the problem of sampling from an unknown probability measure $\nu({\rm dx})$ on $\mathbb{R}^{d}$ for which we only have access to a finite set of training samples $x^{(j)}\sim\nu$, $j=1,\ldots,M$. This problem has recently attracted widespread interest in the context of score-generative or diffusion modeling [18, 7, 17, 19, 22]. If the probability measure $\nu({\rm d}x)$ possesses a probability density function $\pi(x)$, then a popular non-parametric approach to generative modeling is to estimate the score function $s_{\theta}(x)\approx\nabla\log\pi(x)$ by minimizing an appropriate loss function such as
(1) $\mathcal{L}(\theta)=\int_{\mathbb{R}^{d}}\pi(x)\|s_{\theta}(x)-\nabla\log\pi(x)\|^{2}{\rm d}x$
@@
(3) $X(n+1)=X(n)+\epsilon s_{\theta}(X(n))+\sqrt{2}\,\Xi(n),\qquad\Xi(n)\sim{\rm N}(0,\epsilon I),$
for $n\geq 0$, where $\epsilon>0$ denotes the step size and $X(n)$ provides the numerical approximation to the solution of (2) at time $\tau_{n}=n\,\epsilon$. The EM algorithm is initialized at one of the training data points; i.e., $X(0)=x^{(j^{\ast})}$ with $j^{\ast}\in\{1,\ldots,M\}$ appropriately chosen, and the resulting discrete trajectory $X(n)$, $n\geq 1$, delivers approximate samples from the target distribution $\nu({\rm d}x)$.
@@
(5) $X(n+1)=m(X(n);\epsilon)+\sqrt{2}\,\Xi(n),\qquad\Xi(n)\sim{\rm N}(0,\epsilon I).$
@@
We emphasize that the discrete-time formulation (5) can be considered even in case the probability measure $\nu({\rm d}x)$ does not possess a probability density function $\pi(x)$ with respect to the Lebesgue measure on $\mathbb{R}^{d}$; e.g., the measure $\nu$ is concentrated on a submanifold $\mathcal{M}\subset\mathbb{R}^{d}$, as long as the conditional expectation value (4) can be defined appropriately. The Schrödinger bridge approximation to $m(x;\epsilon)$ allows for such an extension [5].
While it has been demonstrated in [5] that (5) works well for low-dimensional problems, the required number of training samples, $M$, increases exponentially in the dimension, $d$, of the samples [21]. In order to remedy this manifestation of the curse of dimensionality, we propose to utilize conditional independence in order to replace the Schrödinger bridge estimator for the conditional expectation value $m(x;\epsilon)\in\mathbb{R}^{d}$ by appropriately localized Schrödinger bridge estimators in each of the $d$ components of $m(x;\epsilon)$. While the proposed localization strategy resembles localization strategies used in the ensemble Kalman filter (EnKF) [3, 16, 2], it is fundamentally different in at least two ways: (i) For Gaussian measures with covariance matrix $\Sigma$, the EnKF would localize the empirical estimator of $\Sigma$ while our approach relies on a localization of $\Sigma^{-1}$ as dictated by conditional independence. (ii) Localized Schrödinger bridge estimators are not restricted to Gaussian measures as long as conditional independence can be established.
@@
While (5) can be used in the general context of score-generative or diffusion modeling, our main motivation is in Bayesian inference and conditional sampling with applications to multi-scale processes. Applications to Bayesian inference, for which $\pi(x)$ takes the role of the prior for given likelihood function $\pi(y|x)$, immediately suggest the modified update
(6) $X(n+1)=m(X(n);\epsilon)-\epsilon\nabla\log\pi(y|X(n))+\sqrt{2}\,\Xi(n),\qquad\Xi(n)\sim{\rm N}(0,\epsilon I).$
Furthermore, a particular choice of $\pi(y|x)$ can be used for conditional sampling [5].
@@
The paper is organized as follows. The Schrödinger bridge formulation for $m(x;\epsilon)$ in (5) is summarized in the subsequent Section 2. The localized variant is developed in Section 3 first for a Gaussian distribution for which $\Sigma^{-1}$ has a tri-diagonal structure and then for general target measure $\nu({\rm d}x)$ for which conditional independence holds. An algorithmic summary is provided in Algorithm 1 and a discussion of numerical properties is provided in Section 3.3. As an application, we consider conditional sampling for a closure problem arising from the multi-scale Lorenz-96 model [9] in Section 4. The paper closes with some conclusions and suggestions for further work.
2. Schrödinger bridge sampler
@@
In this section, we briefly recall how to approximate the conditional estimate (4) using Schrödinger bridges. One first introduces the symmetric matrix $T\in\mathbb{R}^{M\times M}$ of (unnormalized) transition probabilities
@@
for $j,k=1,\ldots,M$. See [5] for a more general definition involving a scaling matrix $K(x)$.
@@
Here $D(v)\in\mathbb{R}^{M\times M}$ denotes the diagonal matrix with diagonal entries provided by $v\in\mathbb{R}^{M}$. We remark that the standard scaling used in Schrödinger bridges would lead to a bistochastic matrix $\tilde{P}$, which is related to (8) by $\tilde{P}=M^{-1}P$.
The next step is to extend the discrete Markov chain (8) to all $x\in\mathbb{R}^{d}$. For that purpose one introduces the vector-valued functions $t(x)\in\mathbb{R}^{M}$ with entries
(10) $t^{(j)}(x)=\exp\left(-\frac{1}{4\epsilon}\|x^{(j)}-x\|^{2}\right)$
@@
which provides a finite-dimensional approximation of the conditional expectation value $\mu(x;\epsilon)$ of the true underlying diffusion process (2).
@@
(14a) $X(n+1/2)$ $=X(n)+\sqrt{2}\,\Xi(n),\qquad\Xi(n)\sim{\rm N}(0,\epsilon I),$
(14b) $X(n+1)$ $=\mathcal{X}\,w(X(n+1/2)),$
@@
which can be viewed as sequential noising and denoising steps. The key property of the Schrödinger bridge sampler is that the final step of the Langevin sampler (14)b amounts to a projection into the convex hull of the samples, independent of the outcome of the noising step (14)a. This renders the sampling scheme numerically stable for any finite sample size $M$. This is in contrast to traditional Langevin samplers such as score generative models which directly solve the typically stiff Langevin equation (2); e.g., in case the probability measure $\nu({\rm d}x)$ is concentrated on a submanifold $\mathcal{M}\subset\mathbb{R}^{d}$, simulating the Langevin equation necessitates computationally costly sufficiently small time steps to resolve the fast attraction toward the submanifold.

### v3

@@ Abstract.
We consider the problem of sampling from an unknown distribution for which only a sufficiently large number of training samples are available. In this paper, we build on previous work combining Schrödinger bridges and plug & play Langevin samplers. A key bottleneck of these approaches is the exponential dependence of the required training samples on the dimension, $d$, of the ambient state space. We propose a localization strategy which exploits conditional independence of conditional expectation values. Localization thus replaces a single high-dimensional Schrödinger bridge problem by $d$ low-dimensional Schrödinger bridge problems over the available training samples. In this context, a connection to multi-head self attention transformer architectures is established. As for the original Schrödinger bridge sampling approach, the localized sampler is stable and geometric ergodic. The sampler also naturally extends to conditional sampling and to Bayesian inference. We demonstrate the performance of our proposed scheme through experiments on a high-dimensional Gaussian problem, on a temporal stochastic process, and on a stochastic subgrid-scale parametrization conditional sampling problem. We also extend the idea of localization to plug & play Langevin samplers using kernel-based denoising in combination with Tweedie’s formula.
Keywords: generative modeling, Langevin dynamics, denoising, Schrödinger bridges, conditional independence, localization, Bayesian inference, conditional sampling, multi-scale closure
@@ 1. Introduction
In this paper, we consider the problem of sampling from an unknown probability measure $\nu({\rm dx})$ on $\mathbb{R}^{d}$ for which we only have access to a finite set of training samples $x^{(j)}\sim\nu$, $j=1,\ldots,M$. This problem has recently attracted widespread interest in the context of score-generative or diffusion modeling [28, 12, 27, 29, 36]. If the probability measure $\nu({\rm d}x)$ possesses a probability density function $\pi(x)$, then a popular non-parametric approach to generative modeling is to estimate the score function $s(x;\theta)\approx\nabla\log\pi(x)$ by minimizing an appropriate loss function such as
(1) $\mathcal{L}(\theta)=\int_{\mathbb{R}^{d}}\|s(x;\theta)-\nabla\log\pi(x)\|^{2}\pi(x){\rm d}x$
@@
(3) $X(n+1)=X(n)+\epsilon s(X(n);\theta)+\sqrt{2\epsilon}\,\Xi(n),\qquad\Xi(n)\sim{\rm N}(0,I),$
for $n\geq 0$, where $\epsilon>0$ denotes the step size and $X(n)$ provides the numerical approximation to the solution of (2) at time $\tau_{n}=n\,\epsilon$. The EM algorithm is initialized at one of the training data points; i.e., $X(0)=x^{(j^{\ast})}$ with $j^{\ast}\in\{1,\ldots,M\}$ appropriately chosen, and the resulting discrete trajectory $X(n)$, $n\geq 1$, delivers approximate samples from the target distribution $\pi(x)$.
@@
(5) $X(n+1)=m(X(n);\epsilon)+\sqrt{S(X(n);\epsilon)}\,\Xi(n),\qquad\Xi(n)\sim{\rm N}(0,I),$
with appropriately defined diffusion matrix $S(x;\epsilon)\in\mathbb{R}^{d\times d}$. Broadly speaking, $m(x;\epsilon)$ controls the drift while $S(x;\epsilon)$ moderates the noise. An obvious choice for $S(x;\epsilon)$ is $S(x;\epsilon)=2\epsilon I$, which corresponds to the EM discretization (3). A data-aware $S(x;\epsilon)$ has been introduced in [10], which is defined as the Schrödinger bridge approximation to the covariance matrix
(6) $\Sigma(x;\epsilon):=\mathbb{E}[X(\epsilon)X(\epsilon)^{\top}|X(0)=x]-\mu(x;\epsilon)\mu(x;\epsilon)^{\top}.$
Provided the measure $\nu({\rm d}x)$ possesses a smooth density $\pi(x)$, it holds asymptotically that
(7) $\Sigma(x;\epsilon)=2\epsilon I+\mathcal{O}(\epsilon^{2}).$
@@
We emphasize that the discrete-time formulation (5) can be considered even in case the probability measure $\nu({\rm d}x)$ does not possess a probability density function $\pi(x)$ with respect to the Lebesgue measure on $\mathbb{R}^{d}$; e.g., the measure $\nu$ is concentrated on a submanifold $\mathcal{M}\subset\mathbb{R}^{d}$, as long as the conditional expectation values (4) and (6) can be defined appropriately. The Schrödinger bridge approximation allows for such an extension [10]. Indeed, the, so called, manifold hypothesis states that many applications of generative modeling lead to measures $\nu({\rm d}x)$ which concentrate on a low-dimensional manifold $\mathcal{M}$ in a high-dimensional ambient space $\mathbb{R}^{d}$ [8, 27, 33].
We note that (5) is closely related to the plug & play unadjusted Langevin sampler (PnP-ULA) of [14], where a denoiser $D(x;\epsilon)$ takes the role of $m(x;\epsilon)$ in (5), which should satisfy
(8) $\nabla\log\pi(x)\approx\frac{D(x;\epsilon)-x}{\epsilon}$
and $S(x;\epsilon)=2\epsilon I$. An additional stabilizing term of the form
(9) $\frac{\epsilon}{\lambda}(P_{\mathcal{C}}(X(n))-X(n))$
is required for the associated PnP-ULA scheme to satisfy an appropriate growth condition. Here $\lambda>0$ is a suitable parameter and $P_{\mathcal{C}}(x)$ projects $x$ onto a compact set $\mathcal{C}\subset\mathbb{R}^{d}$ which should contain most of the probability mass of $\nu({\rm d}x)$. In particular, if $\nu({\rm d}x)$ is supported on a manifold $\mathcal{M}\subset\mathbb{R}^{d}$, then $\mathcal{C}\subseteq\mathcal{M}$. See [14] for more details and [5] for a related approach using a Moreau–Yoshida regularised score function. We note that the Schrödinger bridge sampler (5) has been shown to be stable and geometric ergodic [10] without any additional stabilization term.
@@
The sampler (5) can be used in the general context of score-generative or diffusion modeling, however, our main motivation is in Bayesian inference and in conditional sampling with applications to multi-scale processes. Applications to Bayesian inference, for which $\nu({\rm d}x)$ takes the role of the prior for given likelihood function $\pi(y|x)$, immediately suggest the modified update
(10) $X(n+1)=m(X(n);\epsilon)+\epsilon\nabla\log\pi(y|X(n))+\sqrt{S(X(n);\epsilon)}\,\Xi(n),\qquad\Xi(n)\sim{\rm N}(0,I).$
Furthermore, a particular choice of $\pi(y|x)$ can be used for conditional sampling [10]. We note that (10) fits into the general plug & play approach to data-aware Bayesian inference [31, 3, 14].
While it has been demonstrated in [10] that (5) and (10) work well for low-dimensional problems, the required number of training samples, $M$, increases exponentially in the dimension, $d$, of the samples [35]. In order to remedy this manifestation of the curse of dimensionality, we propose to utilize conditional independence in order to replace the Schrödinger bridge estimator for the conditional expectation value $m(x;\epsilon)\in\mathbb{R}^{d}$ by appropriately localized Schrödinger bridge estimators in each of the $d$ components of $m(x;\epsilon)$ and similarly for the diffusion matrix $S(x;\epsilon)$. The proposed localization strategy resembles localization strategies used in the ensemble Kalman filter (EnKF) [7, 25, 4]; but is fundamentally different in at least two ways: (i) For Gaussian measures with covariance matrix $C$, the EnKF would localize the empirical estimator of $C$ while our approach relies on the localization of the precision matrix $C^{-1}$ as dictated by conditional independence. (ii) Localized Schrödinger bridge estimators are not restricted to Gaussian measures as long as conditional independence can be established. Furthermore, we extend the proposed localization strategy to conditional mean estimators based on kernel denoising [20].
@@
The paper is organized as follows. The Schrödinger bridge formulation for $m(x;\epsilon)$ and $S(x;\epsilon)$ in (5) is summarized in the subsequent Section 2. There we also discuss connections to minimum mean square error (MMSE) denoising and kernel-based denoising, in particular to [20]. The localized variant is subsequently developed in Section 3 first for a Gaussian distribution for which $C^{-1}$ has a tri-diagonal structure and then for general target measure $\nu({\rm d}x)$ for which conditional independence holds. An algorithmic summary is provided in Algorithm 1 and a discussion of numerical properties is provided in Section 3.3. Localization is extended to kernel-based denoising in Section 3.4. We discuss a connection between the Schrödinger bridge sampler and self attention transformers [30] in Remark 2 and its localized variant in the context of multi-head transformers in Remark 3. As applications, we consider sampling temporal stochastic processes in Section 4 and conditional sampling for a closure problem arising from the multi-scale Lorenz-96 model [16] in Section 5. The paper closes with some conclusions and suggestions for further work.
2. Plug & play Langevin sampler
In this section, we summarize two particular variants of plug & play Langevin samplers [14]. The first sampler has been proposed in [10] and is based on a Schrödinger bridge approximation of the Langevin semi-group with invariant measure $\nu({\rm d}x)$ [35]. The second sampler builds upon kernel denoising [20] and Tweedie’s formula [6].
2.1. Schrödinger bridge sampler
@@
In this subsection, we briefly recall how to approximate the conditional estimates (4) and (6) using Schrödinger bridges. One first introduces the symmetric matrix $T\in\mathbb{R}^{M\times M}$ of (unnormalized) transition probabilities
@@
for $j,k=1,\ldots,M$. See [10] for a more general definition involving a state-dependent scaling matrix $K(x)$ and variable bandwidth implementation $K(x)=\rho(x)I$ with $\rho(x)>0$ a suitable scaling function.
@@
Here $D(v)\in\mathbb{R}^{M\times M}$ denotes the diagonal matrix with diagonal entries provided by $v\in\mathbb{R}^{M}$. We remark that the standard scaling used in Schrödinger bridges would lead to a bistochastic matrix $\tilde{P}$, which is related to (12) by $\tilde{P}=M^{-1}P$ [22].
The next step is to extend the discrete Markov chain (12) to all $x\in\mathbb{R}^{d}$. For that purpose one introduces the vector-valued function $t(x)\in\mathbb{R}^{M}$ with entries
(14) $t^{(j)}(x)=\exp\left(-\frac{1}{4\epsilon}\|x-x^{(j)}\|^{2}\right)$
@@
which provides a finite-dimensional approximation of the conditional expectation value $\mu(x;\epsilon)$ of the true underlying diffusion process (2). Note that the conditional mean $m(x;\epsilon)$ lies in the convex hull of the data since $0\leq w(x)\leq 1$ is a probability vector for all $x$
We also recall a data-aware choice of the covariance matrix $S(x;\epsilon)$ [10]. Using (15) and (16), one can define the conditional covariance matrix
(18) $S(x;\epsilon)=\mathcal{X}\,D(w(x))\,\mathcal{X}^{\top}-m(x;\epsilon)\,m(x;\epsilon)^{\top}\in\mathbb{R}^{d\times d},$
which is the empirical covariance matrix associated with the probability vector $w(x)$; compare (6).
@@
(19a) $X(n+1/2)$ $=X(n)+\sqrt{S(X(n);\epsilon)}\,\Xi(n),\qquad\Xi(n)\sim{\rm N}(0,I),$
(19b) $X(n+1)$ $=m(X(n+1/2);\epsilon),$
@@
which can be viewed as sequential noising and denoising steps. The key property of the Schrödinger bridge sampler is that the final step of the Langevin sampler (19b) amounts to a projection into the convex hull of the samples, independent of the outcome of the noising step (19a). This renders the sampling scheme numerically stable for any finite sample size $M$. This is in contrast to traditional Langevin samplers such as score generative models which directly solve the typically stiff Langevin equation (2); e.g., in case the probability measure $\nu({\rm d}x)$ concentrates on a submanifold $\mathcal{M}\subset\mathbb{R}^{d}$, simulating the Langevin equation necessitates computationally costly sufficiently small time steps to resolve the fast attraction toward the submanifold [27].
Remark 2.
We point to a connection of (19b) to self-attention transformer architectures [30]. We recall that the attention function acts on a matrix $Q\in\mathbb{R}^{N\times d}$ of $N$ queries, a matrix $K\in\mathbb{R}^{M\times d}$ of $M$ keys, and a matrix $V\in\mathbb{R}^{M\times d}$ of $M$ values in the form of
(20) $\mbox{Attention}\,(Q,K,V)=\mbox{softmax}\left(\frac{QK^{\rm T}}{\sqrt{d}}\right)V.$
In the context of (19b) we find that $Q=X(n+1/2)^{\rm T}$ and $K=V=\mathcal{X}^{\rm T}$. Hence $N=1$ and the output of the softmax function becomes a probability vector of dimension $1\times M$ which we denote by $\breve{w}$. Note that $\breve{w}\in\mathbb{R}^{1\times M}$ multiplies $V=\mathcal{X}^{\rm T}\in\mathbb{R}^{M\times d}$ from the left resulting in essentially the transpose of what has been used in (17) with, however, a differently defined probability vector. Indeed, the Schrödinger bridge sampler defines the probability vector (15) in a manner closely related to what has been proposed as Sinkformer in [26] and the scaling factor $\sqrt{d}$ in (20) is substituted by $2\epsilon$. More specifically, we note that (11) could be replaced by
(21) $(T)_{jk}=\exp\left(\frac{(x^{(k)})^{\top}x^{(j)}}{2\epsilon}\right)$
without changing the resulting Schrödinger bridge approximation $P$ (albeit with a different scaling vector $v$ compared to (12)). One would then also have to replace (14) and (15) by
(22) $\hat{t}^{(j)}(x)=v^{(j)}\exp\left(\frac{x^{\top}x^{(j)}}{2\epsilon}\right)=\exp\left(\frac{x^{\top}x^{(j)}}{2\epsilon}+\log v^{(j)}\right)$
and
(23) $\hat{w}^{(j)}(x)=\frac{\hat{t}^{(j)}(x)}{\sum_{j=1}^{M}\hat{t}^{(j)}(x)}$
for $j=1,\ldots,M$, respectively, in line with self-attention transformer architectures which do not involve the shift by $\log v^{(j)}$ in (22).
The denoising step (19b) has a gradient structure since
(24) $m(x;\epsilon)=x+\epsilon\nabla\log p_{\epsilon}(x)$
with (unnormalised) density
(25) $p_{\epsilon}(x)=\left(v^{\top}t(x)\right)^{2}$
and
(26) $\nabla\log p_{\epsilon}(x)=-\frac{1}{\epsilon}\frac{\sum_{j=1}^{M}(x-x^{(j)})v^{(j)}t^{(j)}(x)}{v^{\rm T}t(x)}=\frac{1}{\epsilon}(\mathcal{X}\,w(x)-x).$
Hence the proposed sampler can be viewed as an EM approximation of the modified Langevin dynamics
(27) $\dot{X}(\tau)=\nabla\log p_{\epsilon}(X(\tau))+\sqrt{2}\,\dot{W}(\tau).$
A modified score has also been considered in the form of Moreau–Yosida regularised score functions in [5] and smoothed score functions in the form of plug & play priors in [14]. Contrary to those approaches, the modified score $\nabla\log p_{\epsilon}(x)$ arises from the Schrödinger bridge approximation of the semi-group $\exp(\epsilon\mathcal{L})$ with generator $\mathcal{L}$ [21] given by
(28) $\mathcal{L}f=\nabla\log\pi(x)\cdot\nabla f+\Delta f.$
2.2. Kernel denoising and Tweedie’s formula
We note that (19b) is related to MMSE denoising as widely used to reduce random fluctuations in a signal. The connection between score estimation, autoencoders, and denoising has been discussed in [32, 1]. See also the recent survey [20]. However, while MMSE denoising typically considers conditional mean estimators in pseudo-linear form [20] or in the form of auto-encoders [1], our approach relies on (nonlinear) conditional mean estimators of the form (17), which also arise from kernel denoising [20], which is closely related to Tweedie’s formula [6] as we explain next.
Given the data distribution $\pi$ and a scale parameter $\gamma>0$, consider the extended (unnormalized) distribution $\Pi_{\gamma}$ in $(x,x^{\prime})\in\mathbb{R}^{2d}$ defined by
(29) $\Pi_{\gamma}(x,x^{\prime})=\exp\left(-\frac{1}{2\gamma}\|x-x^{\prime}\|^{2}\right)\,\pi(x^{\prime})$
and its (unnormalized) marginal distribution
(30) $\pi_{\gamma}(x)=\int\Pi_{\gamma}(x,x^{\prime})\,{\rm d}x^{\prime}.$
Tweedie’s formula [6] states that
(31) $\nabla\log\pi_{\gamma}(x)=-\frac{1}{\gamma}\left(x-\mathbb{E}[x^{\prime}|x]\right)$
with the conditional expectation value defined by
(32) $\mathbb{E}[x^{\prime}|x]=\frac{\int x^{\prime}\,\Pi_{\gamma}(x,x^{\prime})\,{\rm d}x^{\prime}}{\pi_{\gamma}(x)}.$
We may now replace the data distribution $\pi_{\gamma}$ by the empirical measure over the training samples $\{x^{(j)}\}_{j=1}^{M}$ to obtain the equally weighted (unnormalized) Gaussian kernel density estimator (KDE)
(33) $\tilde{\pi}_{\gamma}(x)=\sum_{j=1}^{M}\exp\left(-\frac{1}{2\gamma}\|x-x^{(j)}\|^{2}\right),$
which, according to (31), leads to the score function
(34) $s(x;\gamma)=\nabla\log\tilde{\pi}_{\gamma}(x)=-\frac{1}{\gamma}\left(x-\frac{\sum_{j=1}^{M}x^{(j)}\exp\left(-\frac{1}{2\gamma}\|x-x^{(j)}\|^{2}\right)}{\sum_{j=1}^{M}\exp\left(-\frac{1}{2\gamma}\|x-x^{(j)}\|^{2}\right)}\right).$
Using this score function in (3) with $\gamma=\epsilon$ results in a scheme of the form (5) with $S(x;\epsilon)=2\epsilon I$ and the conditional mean estimator $m(x;\epsilon)$ being replaced by the denoiser
(35) $D(x;\epsilon):=\frac{\sum_{j=1}^{M}x^{(j)}\exp\left(-\frac{1}{2\epsilon}\|x-x^{(j)}\|^{2}\right)}{\sum_{j=1}^{M}\exp\left(-\frac{1}{2\epsilon}\|x-x^{(j)}\|^{2}\right)}=\mathcal{X}\,\tilde{w}(x),$
where the weight vector $\tilde{w}(x)\in\mathbb{R}^{M}$ is now defined by
(36) $\tilde{w}^{(j)}(x)=\frac{\exp\left(-\frac{1}{2\epsilon}\|x-x^{(j)}\|^{2}\right)}{\sum_{j=1}^{M}\exp\left(-\frac{1}{2\epsilon}\|x-x^{(j)}\|^{2}\right)},\qquad j=1,\ldots,M.$
We find that $m(x;\epsilon)$ and $D(x;\epsilon)$ differ through the additional Sinkhorn weight vector $v\in\mathbb{R}^{M}$ in (15) and the scale parameter $2\gamma=2\epsilon$ in (33) compared to the scale parameter $4\epsilon$ used in (14). The connections drawn in Remark 2 to transformer architectures apply equally to (35).
The results of [35] suggest that (17) provides a more accurate approximation to the conditional expectation value (4) than (35), which is based on the equally weighted Gaussian mixture approximation (33). In particular, it holds that
(37) $m(x;\epsilon):=\exp(\epsilon\mathcal{L})\,{\rm id}(x)=x+\epsilon\nabla\log\pi(x)+\mathcal{O}(\epsilon^{2})$
in the limit $M\to\infty$, while Tweedie’s formula formally leads to
(38) $D(x;\epsilon):=x+\epsilon\nabla\log\tilde{\pi}_{\epsilon}(x)=x+\epsilon\nabla\log\pi(x)+\mathcal{O}(\epsilon^{2}).$
Here ${\rm id}(x)=x$ denotes the identity map. Hence, to leading order in $\epsilon$, both approaches agree. However, while Tweedie’s formula leads to an approximation error that arises from replacing $\pi$ by a regularized density ${\tilde{\pi}}_{\epsilon}$, the Schrödinger bridge sampler leads to higher-order corrections which are consistent with the actual underlying Langevin dynamics. This becomes particularly appealing when implemented together with the data-aware covariance matrix (18) instead of a constant $S(x;\epsilon)=2\epsilon I$ in (5) or when a variable bandwidth is implemented in (11) as was done in [10]. A precise statement will be the subject of future research. We also stress that the Schrödinger bridge sampler can easily be extended to Langevin dynamics with multiplicative noise [10] while such an extension is unclear when based on a KDE.